please find the value of NAOH added , as well as the molarity NAOH
Volume NaOH added M Molarity NAOH
REPORT FORM molar mass of KHC H O4 g/mole Sample # 22,391 22.4315 2222 MAW 21.137 2614 22 2327361 1 3 Masses of KHC H,O4 Samples Mass of bottle + sample 2223 Mass of bottle + remaining material 20.744820.370 04S 0.50 0.11 036 S5.44 $S4CTAS 1790 135 O,00 0.00 0,504 Mass of sample Buret Readings mL Final reading 4,8 mL Initial reading le NaOH
Volume NaOH added M Molarity NAOH
Answer #1

You are determining the molar mass of molecule KHC8H4O4

And to calculate the molar mass of this unknown compound you have to first determine the Molarity of KHC8H4O4

And in quantitative analysis/titration process - to determine the concentration of unknown compound the Molarity of one compound is known which is going to be filled in burette. That means the standard solution of NaOh is prepared in lab whose Molarity should be known to you. So please discuss with your labmates to check the recorded data of concentration of NaOH.

And the volume of NaOh can be calculated by calculating = final concentration - initial concentration

Trial 1 = 55.44 - 4.8 = 50.64 mL

Trial 2 = 7.50 - 0.0 = 7.50 mL

Trial 3 = 17.90 - 0.00 = 17.90 mL

Trial 4 = 7.35 - 0.5049 = 6.8451 mL